A Complete Bibliography of Publications in the Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches

sur les pêcheries du Canada (1940–1949)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.02

Title word cross-reference

2 [102].

Abdominal [85]. Abnormal [102]. Abundance [24, 8]. Acid [20, 152, 26]. Acids [64]. Action [113, 30, 61, 79, 92, 134]. Activities [63]. Activity [48, 62]. Adsorption [65]. Affecting [91, 95]. Agents [103]. Agonid [56]. along [143]. americanus [78]. Amines [93, 94, 99]. Amino [152]. Ammonium [73]. Analyses [36]. Analysis [142]. Animal [142]. Animals [122, 60]. Antibiotics [134]. Antioxidants [126, 153]. approaches [41]. Aquatic [122]. Areas [109]. Artificial [132, 57]. Artificiel [132]. Associated [140]. Atlantic [47, 85, 45, 125, 138, 57]. Autumn [11].

B [120]. B-Vitamins [120]. B.C. [46, 68, 77, 135]. Bacteria [80, 110, 140, 154, 44, 53, 54, 12, 58, 30, 61, 79, 83, 134, 74]. Bacterial [141, 148, 66]. Bactéries [12]. Bacteriological [155, 157]. Bacteriostatic [92]. Bacterium [2]. Baffin [101]. Bases [87]. Basis [143]. Bay [107, 67, 3]. Bear [130]. Beds [57]. Behaviour [77, 29]. Bermuda [139]. Between [70]. Biological [67, 115]. Birds [24]. Blister [111]. Blood [70]. Body [43]. Bottom [68]. Breakdown [103]. British [50, 72, 36]. Brook [18, 49]. Brown [32, 46]. Buffering [21].

Caeca [97]. Canada [5, 151]. Canadian [47, 138, 108, 111, 131, 9, 85]. Canned [73, 124, 36]. Capacity [21]. Capelan [132]. Capelin [132]. Carbonyl [153]. Catch [8]. Catchability [86]. Centra [37]. Certain [23, 64]. Changes [147, 69, 21]. Characteristics [81, 77]. Chemical [83, 123, 136, 126]. Choline [113]. Chromatographic [142]. Ciferri [32]. Circular [146]. Circulation [156]. Cladoceran [14]. Clam [133]. Clams [36]. Clostridium [119, 118]. Clupea [4, 37]. CO [102]. Coast [143, 47, 85]. Coastal [101]. Cod [6, 140, 141, 144, 154, 145, 20, 21, 99, 125, 147, 11]. Coho [71]. coli [27]. Colorimetric [100, 93]. Columbia [72, 50, 36]. Column [4]. Comme [103]. Commensal [133]. Commercial [23, 58]. Comparative [15, 127]. Composition [152]. Compounds [113, 134]. Concept [59]. Condition [28]. Conditions [67]. Conrad [69]. Constant [146]. Constituents [91]. Consumption [33]. containing [7]. Content [70]. Control [126, 136, 153]. Copepods [39]. Corrosion [124, 123]. Counts [141, 97]. Cowichan [46, 68]. Crabs [36]. Creek [77, 135]. Critique [143]. Cruising [129]. Crystal [73]. Cultures [128]. Cultus [34]. Current [146, 112, 106]. currents [41]. Cutthroat [46]. Cycle [70]. Cycles [138]. Cyclical [24].

D [63, 43]. Death [51, 17]. Décomposition [103]. Dehydrated [87]. Density [84]. Depleted [109]. Deposition [91, 76]. Determination [93, 100]. Develop [121]. Development [4, 102, 12, 69, 94]. Developmental [9]. Développement [12]. Differentiation [128]. Dimethylamine [6]. Dimorphism [78]. Disinfectants [54]. Disinfection [123]. Distribution [116, 5, 10]. Diurnal [62, 71]. Dogfish [89]. Drip [42]. Dropping [82]. Drying [132, 96]. Dun [31]. During [155].

Eastern [5]. Eberthella [27]. Eel [146]. Effect [110, 119, 144, 158, 48, 34, 39, 102, 104, 90, 42, 60, 53, 54]. Effectiveness [7]. Effects [82, 156, 52]. Efficiency [135]. Effort [8]. Egg [9]. Egg-Laying [9]. Eggs [109, 102]. Electrodes [117]. Embryonic [9]. emend [32]. Enediols [153]. Enterobacteriaceae [113, 74]. Environment [53]. Enzymatic [158]. Enzyme [95]. Enzymes [23]. epizoum [31]. Escherichia [27]. Esox [131]. Essential [64]. Esterified [63]. Estimation [124]. Estuary [16]. Étude [149]. Eviscerated [99]. Experiment [121]. Experimental [60].

Experiments [72, 61]. Exploitation [8].

Factors [91, 95, 10]. Fall [66]. Fate [22]. Fatness [104]. Fatty [64]. Fauna [68]. Feeding [62]. Female [85]. Fertility [137]. Fertilization [137]. Fillets [140, 141, 144, 154, 145, 147, 15]. Fish [64, 6, 115, 120, 116, 119, 118, 20, 21, 93, 94, 99, 148, 155, 34, 103, 27, 31, 32, 44, 52, 53, 54, 56, 96, 91, 12, 33, 58, 65, 7, 26, 42, 66, 83, 87, 126, 134, 136, 127, 157, 153, 59]. Fishery [152]. Fishes [108, 23, 90]. Flesh [115, 120, 116, 126, 134, 136, 157, 153]. Fluctuations [71]. fontinalis [48, 18, 29, 49]. Food [46, 39, 13, 114]. Formation [148, 73]. Four [131]. Frais [132, 12]. Free [63]. French [149, 103, 132, 12]. Fresh [132, 15, 12]. Fry [77]. Fuca [41]. Fundy [107]. Further [155, 61].

gairdneri [81]. Genus [32]. Germicidal [127]. Gill [97]. Girard [13]. Glucose [103, 58]. Glycerol [89]. Gmelin [76]. Goldfish [129]. Gonad [69]. Goniaulax [150]. gorbuscha [135]. Great [130]. Greenland [106]. Growth [120, 110, 119, 118, 53, 156, 35, 90, 3, 83, 134].

Habits [81]. Halophilic [44, 53, 54, 32]. Hatching [82]. Heat [108]. Heavily [96]. Herring [4, 124, 39, 102, 37, 50, 43, 143]. High [51]. History [18, 55]. Homarus [78]. Hudson [67]. Huile [149]. Hydroxylamine [92].

Ice [25, 7]. Iced [99]. Ices [26, 127]. II [94, 12, 136]. III [99, 153]. Individual [38]. Influence [28]. Inhibiteurs [103]. Inhibiting [103]. Inhibition [83, 95]. Inlet [25, 135]. Intestines [44, 43]. Iodate [103]. Island [101, 143]. IV [27]. IX [64, 21, 53].

Juan [41]. Juice [6, 20, 21].

kennerlyi [121]. Killifish [114].

Lake [84, 130, 34, 130]. Lakes [131]. land [121]. land-locked [121]. Laying [9]. Lepeophtheirus [17, 55]. libellula [101]. Lice [17]. Life [18, 55, 49]. Light [39]. Limiting [10]. Liver [43, 65, 89]. Lobster [73, 70, 78]. Lobsters [86, 9, 85]. locked [121]. Loup [149]. Loup-Marin [149]. lucius [131]. Lunar [139].

Magnesium [73]. Magnesium-Ammonium-Phosphate [73]. Malpeque [3]. Man [22]. Marin [149]. Marine [115, 120, 118, 128, 36, 142]. Maritime [108]. Masset [135]. Matter [142]. Maturity [85]. McClinton [77, 135]. Mean [143]. Measurement [59]. Metabolism [58]. Method [143]. Methods [136]. Micrococcaceae [80]. Microorganisms [125, 87]. Migrating [29]. Migration [98, 88, 71]. Migratory [81]. Molds [32]. Moser [18]. Moult [70]. Mountain [13, 35]. Movement [146]. mucronala [14]. Mud [111]. Mud-Blister [111]. Muscle [6, 20, 21, 93, 94, 99, 70, 42, 66].

N.S. [18]. NaCl [42]. Natural [135]. navalis [3]. nerka [121, 84]. Newfoundland [5]. Nitrates [79]. Nitrite [144, 154, 100, 148, 155, 7, 26, 60, 134]. Nitrite-Reducing [154]. Nitrites [30, 61, 79]. Non [128, 41]. Non-Marine [128]. non-tidal [41]. Northern [131]. Nova [11, 122]. Number [50, 143]. Numbers [14]. Nutritionally [64]. Nutritive [115, 120, 36].

Obituaries [75]. OBITUARY [1, 105]. Observations [2, 130]. Occurrence [151, 16]. Oceanographical [67]. Odontopyxis [56]. Odontosyllis [139]. Oil [149, 43, 89]. Oils [64, 65, 142]. Oncorhynchus [84, 121, 135]. Ostrea [28, 76]. Other [94]. Ova [82]. Oxide [80, 110, 144, 22, 74, 158]. Oyster [76]. Oysters [156, 28, 104, 111, 38].

Pacific [4, 37]. Palatability [147]. pallasii [4, 37]. Paphia [69]. Paralytic [150]. Parr [98]. Passamaquoddy [39]. Patterns [112]. Peculiar [56]. Peptones [157]. Perch [114]. Periodicity [139, 14]. pH [158, 12, 42]. Phosphate [73, 26]. Phosphorus [70]. Physical [77, 136]. Picrate [93]. Pike [131]. Pilchard [43]. Pilchards [72]. Pink [77, 97, 135]. Plankton [16]. Poisoning [150]. Poisson [103, 12]. Polychaeta [107, 67]. Polydora [111]. Population [84]. Populations [66, 143]. Portions [68]. Potassium [103, 103]. Potencies [43]. Predaceous [34, 33]. Preservation [52, 7, 26, 127]. Preservatives [54, 15]. Press [6, 20, 21]. Prevention [123]. Production [6]. Products [115, 120, 152, 36]. Propagation [135]. Prosopium [13]. Proteins [115, 116]. Proximate [36]. Pseudomonas [128, 145]. Pteropods [5]. Purification [149]. putrefaciens [145]. Pycnogonida [47]. Pyloric [97].

Quahaugs [156]. Quality [141, 76].

Rôle [12]. Racial [81]. Rainbow [46]. Rakers [97]. Rancidity [126, 136, 153]. Rate [8]. Rates [9]. Rats [120]. Records [47]. Red [53, 54, 44]. Redaelli [32]. Redds [57]. Reducing [154]. Reduction [80, 144, 158, 148, 155, 74, 34]. Regeneration [45]. Related [113]. Relation [141, 84, 129, 70, 8, 114]. Relative [63]. Relaying [104]. Restocking [109]. Retention [16]. Reversal [38]. Rise [66]. River [46, 68, 18]. Rocky [13, 35]. Role [12]. run [121]. Running [18, 29].

Salinities [25]. Salinity [28]. Salmo [81]. Salmon [45, 146, 34, 84, 109, 121, 62, 24, 51, 98, 88, 71, 77, 97, 135, 43, 33, 17, 57]. salmonicida [2]. salmonis [55]. Salt [93, 27, 31, 32, 44, 52, 53, 54]. Salted [96]. Salting [119]. Salts [52]. Salvelinus [48, 18, 29, 49]. Sarcotaces [151]. Scales [45]. Scapholeberis [14]. Scotia [122]. Scotian [40, 112, 11]. Sea [117, 18, 29, 49, 121, 17]. sea-run [121]. Sea-Running [18, 29]. Sea-water [117]. Seal [149]. Seasonal [69]. Seaward [88]. Seawater [44]. Seaweeds [118]. Séchage [132]. Sex [38, 69]. Sexual [85, 78]. Shallow [25]. Shelf [40, 112]. Shell [76, 133]. Shellfish [150]. Shipworms [3]. Shrimps [36]. Silver [117]. Size [84, 10]. Slime [44]. Smoke [91]. Smoking [119]. Snail [133]. Sockeye [34, 84, 109, 121, 33]. Sodium [148, 7, 26, 60]. Soft [133]. Soft-shell [133]. Sources [64]. Southern [50]. Spawning [11, 57]. Specificity [19]. Speed [129]. Spoilage [6, 140, 20, 21, 155, 83, 59, 99]. Sporendonema [31, 32]. Spring [71]. staminea [69]. Starfish [10]. State [106]. Sterilizing [117]. Strait [41]. Stream [108]. Streams [137]. Stroke [108]. Structure [76]. Studies [6, 20, 21, 155, 27, 31, 32, 44, 52, 53, 54, 31]. Study [149]. Subsequent [82]. Sulpha [134]. Sulphur [116]. Supply [114]. Surface [138, 41, 59]. Survival [119, 34, 27]. Swelling [42]. System [46].

Tagging [72]. tamarensis [150]. Teleostean [82]. Temperature [48, 129, 51, 28]. Temperatures [138, 25]. Tension [102]. Teredo [3]. Themisto [101]. Thiaminase [122]. tidal [41]. Transfer [109]. Transferring [104]. Transports [112]. Trends [138]. Triaenophorus [90]. Triamineoxidease [95, 19]. Trimethylamine [80, 110, 144, 158, 93, 94, 148, 22, 74]. trispinosus [56]. Trout [62, 51, 46, 130, 18, 49]. Tryptic [23]. Tullibee [43]. two [90]. typhosa [27].

Unit [8]. Unsaponifiable [142]. upon [53]. Upstream [71].

V [31]. Value [115, 120, 36, 15, 127, 120]. Vancouver [143]. Variable [88]. Variants [56]. Variation [37]. Variations [62]. Various [52]. versus [24]. Vertebral [4, 37, 50, 143]. VI [32]. Viability [2]. VII [6, 44]. VIII [20, 52]. virginica [28, 76]. Vitamin [63, 43, 65]. Vitamins [120]. Volatile [87, 20].

Water [129, 112, 156, 87, 117]. Waters [101, 40, 5]. West [106, 143]. White [114]. Whitefish [13, 35]. Width [85]. williamsoni [13]. Within [16]. Worm [111].

X [54]. XV [36]. XVI [115]. XVII [120].

Yield [89]. Young [120, 84, 62, 37, 50, 3, 33, 34].

References

Anonymous:1940:O

[1] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5a(1), January 1940. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f40-001.

Duff:1940:OVB

[2] D. C. B. Duff, M. Isobel MacArthur, and Helen G. Thompson. Observations on the viability of Bacterium salmonicida. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5a(1):1–7, January 1940. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f40-002.

Needler:1940:GYS

[3] A. W. H. Needler and A. B. Needler. Growth of young shipworms (Teredo navalis) in Malpeque Bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5a(1):8–10, January 1940. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f40-003.

Gwyn:1940:DVC

[4] Agnes M. Gwyn. The development of the vertebral column of the Pacific herring (Clupea pallasii). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5a(1):11–22, January 1940. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f40-004.

Kerswill:1940:DPW

[5] C. J. Kerswill. The distribution of pteropods in the waters of Eastern Canada and Newfoundland. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5a(1):23–31, January 1940. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f40-005.

Beatty:1940:SFS

[6] S. A. Beatty and V. K. Collins. Studies of fish spoilage: VII. Dimethylamine production in the spoilage of cod muscle press juice. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5a(1):32–35, January 1940. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f40-006.

Tarr:1940:EIC

[7] H. L. A. Tarr and P. A. Sunderland. Effectiveness of ice containing sodium nitrite for fish preservation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5a(1):36–42, January 1940. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f40-007.

Ricker:1940:RCP

[8] W. E. Ricker. Relation of “Catch per Unit Effort” to abundance and rate of exploitation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5a(1): 43–70, January 1940. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f40-008.

Templeman:1940:EDR

[9] W. Templeman. Embryonic developmental rates and egg-laying of Canadian lobsters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5a(1):71–83, January 1940. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f40-009.

Smith:1940:FLD

[10] G. F. M. Smith. Factors limiting distribution and size in the starfish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5a(1):84–103, January 1940. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f40-010.

McKenzie:1940:NSA

[11] R. A. McKenzie. Nova Scotian autumn cod spawning. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5a(2):105–120, February 1940. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f40-011.

Nadeau:1940:PFI

[12] Aristide Nadeau. Le poisson frais: II. Le rôle du pH sur le développement des bactéries. (French) [Fresh fish: II. The role of pH on the development of bacteria]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5a(2):121–130, February 1940. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f40-012.

McHugh:1940:FRM

[13] J. L. McHugh. Food of the Rocky Mountain whitefish Prosopium williamsoni (Girard). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5a(2): 131–137, February 1940. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f40-013.

Smith:1940:PNC

[14] M. W. Smith. Periodicity in the numbers of the cladoceran Scapholeberis mucronala. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5a(2):138–147, February 1940. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f40-014.

Tarr:1940:CVP

[15] H. L. A. Tarr and P. A. Sunderland. The comparative value of preservatives for fresh fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5a(2): 148–163, February 1940. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f40-015.

Rogers:1940:ORP

[16] Harold M. Rogers. Occurrence and retention of plankton within the estuary. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5a(2):164–171, February 1940. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f40-016.

White:1940:SLL

[17] H. C. White. “Sea lice” (Lepeophtheirus) and death of salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5a(2):172–175, February 1940. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f40-017.

White:1940:LHS

[18] H. C. White. Life history of sea-running brook trout (Salvelinus fontinalis) of Moser River, N.S. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5a(2): 176–186, February 1940. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f40-018.

Tarr:1940:ST

[19] H. L. A. Tarr. Specificity of triamineoxidease. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5a(2):187–196, February 1940. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f40-019.

Collins:1941:SFSa

[20] V. K. Collins. Studies of fish spoilage: VIII. Volatile acid of cod muscle press juice. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(3): 197–202, March 1941. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f40-020.

Collins:1941:SFSb

[21] V. K. Collins, C. C. Kuchel, and S. A. Beatty. Studies of fish spoilage: IX. Changes in buffering capacity of cod muscle press juice. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(3):203–210, March 1941. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f40-021.

Tarr:1941:FTO

[22] H. L. A. Tarr. The fate of trimethylamine oxide and trimethylamine in man. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(3):211– 216, March 1941. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f40-022.

Johnston:1941:TEC

[23] W. W. Johnston. Tryptic enzymes from certain commercial fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(3):217–226, March 1941. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f40-023.

Huntsman:1941:CAB

[24] A. G. Huntsman. Cyclical abundance and birds versus salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(3):227–235, March 1941. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f40-024.

Needler:1941:TSU

[25] A. W. H. Needler. Temperatures and salinities under the ice in a shallow inlet. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(3):236– 243, March 1941. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f40-025.

Tarr:1941:ISN

[26] H. L. A. Tarr and P. A. Sunderland. Ices with sodium nitrite and sodium acid phosphate for fish preservation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(3):244–248, March 1941. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f40-026.

Frank:1941:SSFa

[27] M. Frank and E. Hess. Studies on salt fish: IV. Survival of Eberthella typhosa and Escherichia coli on salt fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(3):249–252, March 1941. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f40-027.

Medcof:1941:ITS

[28] J. C. Medcof and A. W. H. Needler. The influence of temperature and salinity on the condition of oysters (Ostrea virginica). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(3):253–257, March 1941. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f40-028.

White:1941:MBS

[29] H. C. White. Migrating behaviour of sea-running Salvelinus fontinalis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(3):258–264, March 1941. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f40-029.

Tarr:1941:ANB

[30] H. L. A. Tarr. The action of nitrites on bacteria. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(3):265–275, March 1941. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f40-030.

Frank:1941:SSFb

[31] M. Frank and E. Hess. Studies on salt fish: V. Studies on Sporendonema epizoum from “Dun” salt fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(3):276–286, March 1941. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f40-031.

Frank:1941:SSFc

[32] M. Frank and E. Hess. Studies on salt fish: VI. Halophilic brown molds of the genus Sporendonema emend. Ciferri et Redaelli. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(3):287–292, March 1941. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f40-032.

Ricker:1941:CYS

[33] W. E. Ricker. The consumption of young sockeye salmon by predaceous fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(3):293– 313, March 1941. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f40-033.

Foerster:1941:ERP

[34] R. E. Foerster and W. E. Ricker. The effect of reduction of predaceous fish on survival of young sockeye salmon at Cultus Lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(4):315–336, April 1941. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f40-034.

McHugh:1941:GRM

[35] T. L. McHugh. Growth of the Rocky Mountain whitefish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(4):337–343, April 1941. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f40-035.

Pugsley:1941:NVM

[36] L. I. Pugsley. The nutritive value of marine products: XV. Proximate analyses of canned British Columbia crabs, shrimps and clams. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(4):344–346, April 1941. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f40-036.

McHugh:1941:VVC

[37] J. L. McHugh. Variation of vertebral centra in young Pacific herring (Clupea pallasii). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(4): 347–360, April 1941. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f40-037.

Needler:1941:SRI

[38] Alfreda Berkeley Needler. Sex reversal in individual oysters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(4):361–364, April 1941. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f40-038.

Johnson:1941:ELC

[39] W. H. Johnson. Effect of light on copepods as food for Passamaquoddy herring. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(4):365– 376, April 1941. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f40-039.

Hachey:1941:WSS

[40] H. B. Hachey. The waters of the Scotian Shelf. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(4):377–397, April 1941. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f40-040.

Tully:1941:SNT

[41] John P. Tully. Surface non-tidal currents in the approaches to Juan de Fuca Strait. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5b(4):398– 409, April 1941. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f40-041.

Tarr:1942:EPN

[42] H. L. A. Tarr. Effect of pH and NaCl on swelling and drip in fish muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5c(5):411–427, May 1942. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f40-042.

Pugsley:1942:VDP

[43] L. I. Pugsley. Vitamin A and D potencies of oil from body, liver and intestines of pilchard, herring, salmon and tullibee. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5c(5):428–437, May 1942. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f40-043.

Hess:1942:SSFa

[44] Ernest Hess. Studies on salt fish: VII. Red halophilic bacteria in seawater and fish slime and intestines. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5c(5):438–439, May 1942. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f40-044.

Blair:1942:RSA

[45] A. A. Blair. Regeneration of the scales of Atlantic salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5c(5):440–447, May 1942. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f40-045.

Idyll:1942:FRC

[46] Clarence Idyll. Food of rainbow, cutthroat and brown trout in the Cowichan River System, B.C. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5c(5):448–458, May 1942. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f40-046.

Giltay:1942:NRP

[47] Louis Giltay. New records of pycnogonida from the Canadian Atlantic Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5c(5):459– 460, May 1942. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f40-047.

Elson:1942:ETA

[48] Paul F. Elson. Effect of temperature on activity of Salvelinus fontinalis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5c(5):461–470, May 1942. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f40-048.

White:1942:SLB

[49] H. C. White. Sea life of the brook trout (Salvelinus fontinalis). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5c(5):471–473, May 1942. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f40-049.

McHugh:1942:VNY

[50] J. L. McHugh. Vertebral number of young herring in Southern British Columbia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5c(5):474– 484, May 1942. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f40-050.

Huntsman:1942:DST

[51] A. G. Huntsman. Death of salmon and trout with high temperature. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 5c(5):485–501, May 1942. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f40-051.

Hess:1942:SSFb

[52] Ernest Hess. Studies on salt fish: VIII. Effects of various salts on preservation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6a(1):1–9, January 1942. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f42-001.

Hess:1942:SSFc

[53] Ernest Hess. Studies on salt fish: IX. Effect of environment upon growth of red halophilic bacteria. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6a(1):10–16, January 1942. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f42-002.

Hess:1942:SSFd

[54] Ernest Hess and N. E. Gibbons. Studies on salt fish: X. Effect of disinfectants and preservatives on red halophilic bacteria. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6a(1):17–23, January 1942. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f42-003.

White:1942:LHL

[55] H. C. White. Life history of Lepeophtheirus salmonis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6a(1):24–29, January 1942. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f42-004.

Hubbs:1942:PVA

[56] Carl L. Hubbs. Peculiar variants of the agonid fish Odontopyxis trispinosus. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6a(1):30–36, January 1942. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f42-005.

White:1942:ASR

[57] H. C. White. Atlantic salmon redds and artificial spawning beds. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6a(1):37–44, January 1942. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f42-006.

Sigurdsson:1942:GMB

[58] G. J. Sigurdsson and A. J. Wood. Glucose metabolism of bacteria from commercial fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6a(1): 45–52, January 1942. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f42-007.

Wood:1942:SCM

[59] A. J. Wood, G. J. Sigurdsson, and W. J. Dyer. The surface concept in measurement of fish spoilage. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6a(1):53–62, January 1942. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f42-008.

Tarr:1942:ESN

[60] H. L. A. Tarr and N. M. Carter. The effect of sodium nitrite on experimental animals. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6a(1): 63–73, January 1942. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f42-009.

Tarr:1942:ANB

[61] H. L. A. Tarr. The action of nitrites on bacteria: Further experiments. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6a(1):74–89, January 1942. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f42-010.

Hoar:1942:DVF

[62] William S. Hoar. Diurnal variations in feeding activity of young salmon and trout. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6a(1):90–101, January 1942. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f42-011.

Bailey:1943:RAF

[63] B. E. Bailey. Relative activities of free and esterified vitamin D. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6b(2):103–108, February 1943. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f42-012.

Bailey:1943:FOI

[64] B. E. Bailey. Fish oils: IX. Certain fish oils as sources of nutritionally essential fatty acids. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6b(2): 109–112, February 1943. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f42-013.

Swain:1943:AVF

[65] Lyle A. Swain. The adsorption of vitamin A from fish liver oils. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6b(2):113–118, February 1943. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f42-014.

Tarr:1943:RFB

[66] H. L. A. Tarr. Rise and fall of bacterial populations in fish muscle. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6b(2):119–128, February 1943. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f42-015.

Berkeley:1943:BOC

[67] E. Berkeley and C. Berkeley. Biological and oceanographical conditions in Hudson Bay: 11. Polychaeta from Hudson Bay. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6b(2):129–132, February 1943. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f42-016.

Idyll:1943:BFP

[68] Clarence Purvis Idyll. Bottom fauna of portions of the Cowichan River, B.C. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6b(2):133–139, February 1943. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f42-017.

Quayle:1943:SGD

[69] D. B. Quayle. Sex, gonad development and seasonal gonad changes in Paphia staminea Conrad. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6b(2):140–151, February 1943. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f42-018.

Hollett:1943:RBM

[70] Andrew Hollett. Relation between moult cycle and phosphorus content of blood and muscle in lobster. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6b(2):152–157, February 1943. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f42-019.

Neave:1943:DFU

[71] Ferris Neave. Diurnal fluctuations in the upstream migration of Coho and spring salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6b(2): 158–163, February 1943. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f42-020.

Hart:1943:TEB

[72] John Lawson Hart. Tagging experiments on British Columbia pilchards. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6b(2):164–182, February 1943. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f42-021.

Hollett:1943:MAP

[73] Andrew Hollett. Magnesium-ammonium-phosphate crystal formation in canned lobster. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6b(2): 183–193, February 1943. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f42-022.

Wood:1943:RTO

[74] A. J. Wood and Elizabeth A. Baird. Reduction of trimethylamine oxide by bacteria: 1. The Enterobacteriaceae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6b(2):194–201, February 1943. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f42-023.

Anonymous:1944:O

[75] Anonymous. Obituaries. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6c(3):203–208, March 1944. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f42-024.

Medcof:1944:SDQ

[76] J. C. Medcof. Structure, deposition and quality of oyster shell (Ostrea virginica Gmelin). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6c(3): 209–216, March 1944. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f42-025.

Pritchard:1944:PCB

[77] A. L. Pritchard. Physical characteristics and behaviour of pink salmon fry at McClinton Creek, B.C. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6c(3):217–227, March 1944. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f42-026.

Templeman:1944:SDL

[78] Wilfred Templeman. Sexual dimorphism in the lobster (Homarus americanus). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6c(3):228– 232, March 1944. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f42-027.

Tarr:1944:ANN

[79] H. L. A. Tarr. Action of nitrates and nitrites on bacteria. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6c(3):233–242, March 1944. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f42-028.

Baird:1944:RTO

[80] Elizabeth A. Baird and A. J. Wood. Reduction of trimethylamine oxide by bacteria:2. The micrococcaceae. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6c(3):243–244, March 1944. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f42-029.

Neave:1944:RCM

[81] Ferris Neave. Racial characteristics and migratory habits in Salmo gairdneri. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6c(3):245–251, March 1944. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f42-030.

Battle:1944:EDS

[82] Helen I. Battle. Effects of dropping on the subsequent hatching of teleostean ova. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6c(3): 252–256, March 1944. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f42-031.

Tarr:1944:CIG

[83] H. L. A. Tarr. Chemical inhibition of growth of fish spoilage bacteria. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6c(3):257–266, March 1944. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f42-032.

Foerster:1944:RLP

[84] R. E. Foerster. The relation of lake population density to size of young sockeye salmon (Oncorhynchus nerka). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6c(3):267–280, March 1944. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f42-033.

Templeman:1944:AWS

[85] Wilfred Templeman. Abdominal width and sexual maturity of female lobsters on Canadian Atlantic Coast. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6c(3):281–290, March 1944. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f42-034.

Smith:1944:CL

[86] G. F. M. Smith. The catchability of lobsters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6c(3):291–301, March 1944. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f42-035.

Tarr:1945:WMV

[87] H. L. A. Tarr. Water, microorganisms and volatile bases in dehydrated fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6d(4):303–310, March 1945. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f42-036.

Huntsman:1945:VSM

[88] A. G. Huntsman. Variable seaward migration of salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6d(4):311–325, March 1945. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f42-037.

Swain:1945:GYD

[89] Lyle A. Swain and B. H. Morton. Glycerol yield from dogfish liver oil. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6d(4):326–333, March 1945. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f42-038.

Miller:1945:ETG

[90] Richard B. Miller. Effect of Triaenophorus on growth of two fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6d(4):334–337, March 1945. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f42-039.

Linton:1945:FAD

[91] E. P. Linton and H. V. French. Factors affecting deposition of smoke constituents on fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6d(4): 338–348, March 1945. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f42-040.

Tarr:1945:BAH

[92] H. L. A. Tarr. Bacteriostatic action of hydroxylamine. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6d(4):349–350, March 1945. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f42-041.

Dyer:1945:AFMa

[93] W. J. Dyer. Amines in fish muscle: I. Colorimetric determination of trimethylamine as the picrate salt. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6d(5):351–358, May 1945. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f42-042.

Dyer:1945:AFMb

[94] W. J. Dyer and Yvonne A. Mounsey. Amines in fish muscle: II. Development of trimethylamine and other amines. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6d(5):359–367, May 1945. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f42-043.

Neilands:1945:FAT

[95] J. B. Neilands. Factors affecting triamineoxidease: I. Inhibition of the enzyme. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6d(5):368– 379, May 1945. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f42-044.

Linton:1945:DHS

[96] E. P. Linton and A. L. Wood. Drying of heavily salted fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6d(5):380–391, May 1945. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f42-045.

Pritchard:1945:CGR

[97] A. L. Pritchard. Counts of gill rakers and pyloric caeca in pink salmon. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6d(5):392–398, May 1945. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f42-046.

Huntsman:1945:MSP

[98] A. G. Huntsman. Migration of salmon Parr. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6d(5):399–402, May 1945. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f42-047.

Dyer:1946:AFM

[99] W. J. Dyer, F. E. Dyer, and M. Snow. Amines in fish muscle: III. Spoilage of iced eviscerated cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6e(6): 403–413, June 1946. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f42-048.

Dyer:1946:CND

[100] W. J. Dyer. Colorimetric nitrite determination. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6e(6):414–418, June 1946. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f42-049.

Dunbar:1946:TLB

[101] M. J. Dunbar. On Themisto libellula in Baffin Island coastal waters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6e(6):419–434, June 1946. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f42-050.

Kelley:1946:EAC

[102] Alice M. Kelley. Effect of abnormal CO_{2} tension on development of herring eggs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6e(6):435– 440, June 1946. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f42-051.

Fougere:1946:LPG

[103] Henri Fougère. L’iodate de potassium et le glucose comme agents inhibiteurs de la décomposition du poisson. (French) [Potassium iodate and glucose as fish breakdown inhibiting agents]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6e(6):441–448, June 1946. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f42-052.

Medcof:1946:ERT

[104] J. C. Medcof. Effect of relaying and transferring on fatness of oysters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6e(6):449–455, June 1946. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f42-053.

Anonymous:1946:O

[105] Anonymous. OBITUARY. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6e(7):457–459, July 1946. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f42-054.

Dunbar:1946:SWG

[106] M. J. Dunbar. The state of the West Greenland Current up to 1944. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6e(7):460–471, July 1946. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f42-055.

Baillie:1946:PBF

[107] W. H. T. Baillie. Polychaeta from the Bay of Fundy 1911–1912. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6e(7):472–475, July 1946. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f42-056.

Huntsman:1946:HSC

[108] A. G. Huntsman. Heat stroke in Canadian Maritime stream fishes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6e(7):476–482, July 1946. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f42-057.

Foerster:1946:RDS

[109] R. E. Foerster. Restocking depleted sockeye salmon areas by transfer of eggs. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6e(7):483– 490, July 1946. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f42-058.

Castell:1946:ETO

[110] Chas. H. Castell. Effect of trimethylamine oxide on the growth of bacteria. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6e(7):491–497, July 1946. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f42-059.

Medcof:1946:MBW

[111] J. C. Medcof. The mud-blister worm, polydora, in Canadian oysters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 6e(7):498–505, July 1946. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f42-060.

Hachey:1947:WTC

[112] H. B. Hachey. Water transports and current patterns for the Scotian Shelf. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7a(1):1–16, January 1947. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f47-001.

Dyer:1947:AEC

[113] Frances E. Dyer and A. J. Wood. Action of Enterobacteriaceae on choline and related compounds. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7a(1):17–21, January 1947. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f47-002.

Smith:1947:FKW

[114] M. W. Smith. Food of killifish and white perch in relation to supply. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7a(1):22–34, January 1947. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f47-003.

Beveridge:1947:NVMa

[115] J. M. R. Beveridge. The nutritive value of marine products: XVI. The biological value of fish flesh proteins. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7a(1):35–49, January 1947. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f47-004.

Beveridge:1947:SDF

[116] J. M. R. Beveridge. Sulphur distribution in fish flesh proteins. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7a(2):51–54, February 1947. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f47-005.

Castell:1947:SES

[117] C. H. Castell, D. G. Ellis, and G. W. Anderson. Silver electrodes for sterilizing sea-water. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7a(2): 55–61, February 1947. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f47-006.

Castell:1947:GCS

[118] Chas. H. Castell. Growth of Clostridium in seaweeds and marine fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7a(2):62–69, February 1947. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f47-007.

Castell:1947:ESS

[119] C. H. Castell and G. W. Anderson. Effect of salting and smoking on survival and growth of Clostridium in fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7a(2):70–73, February 1947. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f47-008.

Beveridge:1947:NVMb

[120] J. M. R. Beveridge. The nutritive value of marine products: XVII. Value of B-vitamins in fish flesh for growth of young rats. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7a(2):74–87, February 1947. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f47-009.

Foerster:1947:EDS

[121] R. E. Foerster. Experiment to develop sea-run from land-locked sockeye salmon (Oncorhynchus nerka kennerlyi). Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7a(2):88–93, February 1947. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f47-010.

Neilands:1947:TAA

[122] J. B. Neilands. Thiaminase in aquatic animals of Nova Scotia. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7a(2):94–99, February 1947. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f47-011.

Tarr:1947:CDC

[123] H. L. A. Tarr. Chemical disinfection and corrosion prevention. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7a(3):101–115, March 1947. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f47-012.

Hollett:1947:ECC

[124] Andrew Hollett. Estimation of corrosion in canned herring. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7a(3):116–127, March 1947. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f47-013.

Dyer:1947:MAC

[125] Frances E. Dyer. Microorganisms from Atlantic cod. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7a(3):128–136, March 1947. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f47-014.

Tarr:1947:CRF

[126] H. L. A. Tarr. Control of rancidity in fish flesh: I. Chemical antioxidants. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7a(3):137–154, March 1947. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f47-015.

Tarr:1948:CVG

[127] H. L. A. Tarr. Comparative value of germicidal ices for fish preservation. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7b(4):155–161, April 1948. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f47-016.

Castell:1948:DMN

[128] C. H. Castell. Differentiation of marine and non-marine cultures of Pseudomonas. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7b(4):162– 168, April 1948. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f47-017.

Fry:1948:CSG

[129] F. E. J. Fry and J. S. Hart. Cruising speed of goldfish in relation to water temperature. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7b(4): 169–175, April 1948. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f47-018.

Miller:1948:OLT

[130] R. B. Miller and W. A. Kennedy. Observations on the lake trout of Great Bear Lake. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7b(4): 176–189, April 1948. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f47-019.

Miller:1948:PEL

[131] R. B. Miller and W. A. Kennedy. Pike (Esox lucius) from four Northern Canadian lakes. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7b(4): 190–199, April 1948. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f47-020.

Fougere:1948:SAD

[132] H. Fougère. Le séchage artificiel du capelan frais. (French) [The artificial drying of fresh capelin]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7b(4):200–211, April 1948. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f47-021.

Medcof:1948:SCS

[133] J. C. Medcof. A snail commensal with the soft-shell clam. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7b(5):219–220, May 1948. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f47-022.

Tarr:1948:ASC

[134] H. L. A. Tarr and Catherine P. Deas. Action of sulpha compounds, antibiotics and nitrite on growth of bacteria in fish flesh. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7b(5):221–223, May 1948. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f47-023.

Pritchard:1948:ENP

[135] A. L. Pritchard. Efficiency of natural propagation of the pink salmon (Oncorhynchus gorbuscha) in McClinton Creek, Masset Inlet, B.C. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7b(5):224–236, May 1948. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f47-024.

Tarr:1948:CRF

[136] H. L. A. Tarr. Control of rancidity in fish flesh: II. Physical and chemical methods. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7b(5):237– 247, May 1948. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f47-025.

Huntsman:1948:FFS

[137] A. G. Huntsman. Fertility and fertilization of streams. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7b(5):248–253, May 1948. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f47-026.

Hachey:1948:TCS

[138] H. B. Hachey and H. J. McLellan. Trends and cycles in surface temperatures of the Canadian Atlantic. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7b(6):355–362, June 1948. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f47-027.

Huntsman:1948:OBL

[139] A. G. Huntsman. Odontosyllis at Bermuda and lunar periodicity. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7b(6):363–369, June 1948. CO-DEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/

doi/full/10.1139/f47-028.

Castell:1948:BAS

[140] C. H. Castell and G. W. Anderson. Bacteria associated with spoilage of cod fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7b(6):370– 377, June 1948. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f47-029.

Castell:1948:RBC

[141] C. H. Castell, G. W. Anderson, and Hilliard Pivnick. Relation of bacterial counts to quality of cod fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7b(6):378–388, June 1948. CODEN JFRBAK. ISSN 0015-296X. URL

https://cdnsciencepub.com/doi/full/10.1139/f47-030.

Swain:1948:CAU

[142] Lyle A. Swain. Chromatographic analysis of the unsaponifiable matter of marine animal oils. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7b(6): 389–402, June 1948. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f47-031.

Tester:1949:PHA

[143] Albert L. Tester. Populations of herring along the West Coast of Vancouver Island on the basis of mean vertebral number, with a critique of the method. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7c(7): 403–420, July 1949. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f47-032.

Castell:1949:ENR

[144] C. H. Castell. Effect of nitrite on reduction of trimethylamine oxide in cod fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7c(7):421– 429, July 1949. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f47-033.

Castell:1949:PPC

[145] C. H. Castell, J. F. Richards, and Isobel Wilmot. Pseudomonas putrefaciens from cod fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7c(7): 430–431, July 1949. CODEN JFRBAK. ISSN 0015-296X. URL https:/

/cdnsciencepub.com/doi/full/10.1139/f47-034.

Davidson:1949:SEM

[146] Viola M. Davidson. Salmon and eel movement in constant circular current. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7c(7):432–448, July 1949. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f47-035.

Dyer:1949:CPC

[147] F. E. Dyer and W. J. Dyer. Changes in the palatability of cod fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7c(8):449–460, August 1949. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f47-036.

Dyer:1949:BRS

[148] W. J. Dyer. Bacterial reduction of sodium nitrite and formation of trimethylamine in fish. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7c(8): 461–470, August 1949. CODEN JFRBAK. ISSN 0015-296X. URL https:

//cdnsciencepub.com/doi/full/10.1139/f47-037.

Dugal:1949:EPL

[149] L.-C. Dugal and A. Cardin. Étude sur la purification de l’huile de loupmarin. (French) [Study on the purification of seal oil]. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7c(8):471–489, August 1949. CODEN JFR-

BAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/full/
10.1139/f47-038.

Needler:1949:PSP

[150] Alfreda Berkeley Needler. Paralytic shellfish poisoning and Goniaulax tamarensis. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7c(8):490– 504, August 1949. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f47-039.

Kuitunen-Ekbaum:1949:OSC

[151] E. Kuitunen-Ekbaum. The occurrence of sarcotaces in Canada. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7c(9):505–512, September 1949. CODEN JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.

com/doi/full/10.1139/f47-040.

Deas:1949:AAC

[152] Catherine P. Deas and H. L. A. Tarr. Amino acid composition of fishery products. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7c(9):513–521, September 1949. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f47-041.

Tarr:1949:CRF

[153] H. L. A. Tarr and N. E. Cooke. Control of rancidity in fish flesh: III. Carbonyl enediols as antioxidants. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7c(9):522–527, September 1949. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f47-042.

Castell:1949:NRB

[154] C. H. Castell. Nitrite-reducing bacteria on cod fillets. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7c(9):528–535, September 1949. CODEN

JFRBAK. ISSN 0015-296X. URL https://cdnsciencepub.com/doi/
full/10.1139/f47-043.

Dyer:1949:FSB

[155] W. J. Dyer and C. H. Castell. Further studies on the bacteriological reduction of nitrite in fish during spoilage. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7c(9):536–544, September 1949. CODEN JFRBAK. ISSN 0015-

296X. URL https://cdnsciencepub.com/doi/full/10.1139/f47-044.

Kerswill:1949:EWC

[156] C. J. Kerswill. Effects of water circulation on the growth of quahaugs and oysters. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7c(9):545–551, September 1949. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f47-045.

Tarr:1949:BPF

[157] H. L. A. Tarr and Catherine P. Deas. Bacteriological peptones from fish flesh. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7c(9):552–560, September 1949. CODEN JFRBAK. ISSN 0015-296X. URL https://

cdnsciencepub.com/doi/full/10.1139/f47-046.

Castell:1949:EPE

[158] C. H. Castell and J. M. Snow. The effect of pH on the enzymatic reduction of trimethylamine oxide. Journal of the Fisheries Research Board of Canada = Journal de l’Office des recherches sur les pêcheries du Canada, 7c(9):561–562, September 1949. CODEN JFRBAK. ISSN 0015-296X.

URL https://cdnsciencepub.com/doi/full/10.1139/f47-047.